OpenStudy (anonymous):

Find the exponential function f(x) = a^x,intersect a point (-2,9)

OpenStudy (anonymous):

plug in x=-2 and f(-2) = 9, then solve for 'a'.

OpenStudy (anonymous):

but i can not get the answer right

OpenStudy (anonymous):

the point (-2,9) is in the form (x,f(x)) if you plug it in \[f(x)=a^x=9=a^{-2}\] You have to use logarithms to solve for a. \[a^x=y\] \[log_a(y)=x\]

OpenStudy (anonymous):

a=1/3?

OpenStudy (anonymous):

answer is (1/3)^x?

OpenStudy (anonymous):

yes